### v4

# Gravitomagnetism from Temporal Dimensional Reduction

Abstract
We dimensionally reduce the Taub-NUT metric to three spatial dimensions by treating time as the extra curled dimension. Since closed timelike curves can exist in the Taub-NUT framework, such a dimensional reduction is justified. Thus, the results end up with $3$-dimensional corresponding Einstein field equations plus corresponding Maxwell type equations for a gravitomagnetic field, associated with the NUT charge, which also acts as a source for the Einstein field equations. In this approach, the Taub-NUT metric can be envisaged as a $(1+3)$-dimensional analogue of the $(1+4)$-dimensional metric of the Kaluza-Klein theory. Hence, in four dimensions, it unifies gravitation and gravitomagnetism, associated with the NUT charge, in the same footing that the Kaluza-Klein theory unifies gravitation and electromagnetism in five dimensions. In fact, in this way, a new gravitomagnetic field emerges from the temporal dimensional reduction. In another word, when the time coordinate is considered curled to a non-observable size, gravitation and gravitomagnetism, associated with the NUT charge, appear as two distinct fields. We also introduce a relation between the $4$-dimensional gravitational constant and the NUT charge.
PACS numbers: 04.20.Gz; 04.20.Cv; 04.20.-q; 04.20.Jb; 04.50.+h
Keywords: Kaluza-Klein Theory; Taub-NUT Metric; Gravitomagnetism; Closed Timelike Curves
1 Introduction
It is known that gravitation and electromagnetism can be unified within the framework of the $(1+4)$-dimensional Kaluza-Klein (KK) theory, which, after compactifying the extra dimension, reduces to the Einstein equations of general relativity (GR) and the Maxwell equations in $(1+3)$-dimensional spacetime through such a purely geometric theory. This theory was proposed in the 1920s by two physicists, Theodor Kaluza and Oskar Klein [1][3]. The main idea behind this theory is to extend the $4$-dimensional spacetime of GR that describes the behavior of gravity, by adding an extra dimension. Kaluza preformed the calculations by taking into account a cylindrical condition, that is, assuming that a fifth dimension exists but no physical quantities depend upon it. Another approach to explaining the apparent $4$-dimensional nature of the universe was introduced by Oskar Klein [2, 3], who hypothesized that the lack of dependence on an extra dimension was because it was compactified to a very small scale, rolled-up in a tinny circle, explaining the fact that such an extra spacial dimension cannot be observed. This idea, commonly known as compactification, proposes a fifth length-like coordinate that has a circular topology and an imperceptible scale [4][12]. This is considered as a type of local gauge symmetry in the associated $4$-dimensional spacetime. Hence, in KK theory, it is assumed that such an extra dimension is responsible for the electromagnetic force, which is not explained by GR. In fact, this theory proposes that the electromagnetic field is not an independent field, but arises from the extra dimension, similar to gravity arises from the curvature of the four dimensions of spacetime.
The KK theory was an important precursor to modern theories of particle physics, such as string theory and supersymmetry [13]. Although this theory has not been fully confirmed by experimental evidence, [^3] Also, in Ref. [14], it has been claimed that the $4$-dimensional spacetime assumed in the KK theory does not exist. it remains an active area of research because it provides a framework for exploring the unification of fundamental forces and the nature of spacetime itself. Since then, it has also become one of the essential foundations for modern theories of particle physics, such as string theory, brane theory, M-theory [15][17], and the theories that led to Yang-Mills theories. In fact, this theory has inspired a great deal of research into the idea of unification of the fundamental forces of nature. However, an ambiguous aspect of the KK theory is the prediction of a dilaton field, which has not yet been observed experimentally.
On the other hand, the Taub-NUT (Newman-Tamburino-Unti) solution [18, 19] to the Einstein equations of GR has been extensively studied in the context of gravitational physics and mathematical relativity [20][22]. One the interesting aspects of this solution is the existence of the NUT charge, which is the source of the gravitomagnetic field [^4] Such a gravitomagnetic field is different from the one resulting from the weak-field and slow-motion limit of GR [23], and was experimentally confirmed in $2011$ [24]. due to the gravitomagnetic monopoles [25][27], the gravitational analogue of the magnetic monopoles [28]. However, a problem with the Taub-NUT solution is the string singularity called the Misner string. One way to eliminate the Misner string is to make the time coordinate periodic [29], which leads to the existence of closed time-like curves (CTCs). CTCs are a hallmark of the Taub-NUT solution and have profound implications for causality in GR. A CTC is a path through spacetime that loops back on itself while remaining time-like throughout. Essentially, it allows for a type of time travel in which the traveler can revisit their past, which violates causality. Causality is a foundational concept in physics, ensuring that in all frames of reference, causes precede effects. The existence of CTCs challenges this principle by allowing situations in which the distinction between cause and effect is broken down. Hence, such a paradoxical behavior implies that the usual interpretation of spacetime as a causal structure, in the presence of CTCs, needs to be modified or scrutinized [30].
However, CTCs are not unique to the Taub-NUT metric, but appear in several solutions of the Einstein field equations, particularly in spacetimes with unusual topology or specific symmetry properties. Notable examples include: The Gödel solution as a rotational cosmological model that admits CTCs [31] and violates global causality everywhere. Also, the interior region (inside the event horizon) of a rotating black hole, as indicated by the Kerr and Kerr-Newman metrics [32], can host CTCs, particularly near the ring singularity. These CTCs are confined to the interior of black holes and do not affect causality in the external spacetime [32]. Traversable wormholes can create CTCs if their ends experience different time dilation effects [33]. Certain configurations of rotating cosmic strings can admit CTCs due to frame-dragging effects [34]. Cylindrical distributions of rotating dust, called the van Stockum dust, can produce CTCs, depending on the density and angular velocity of the dust [35].
The existence of CTCs in various solutions of the Einstein equations has sparked significant debate and research. The causality violations are often considered as artifacts of mathematical structure rather than realizable physical scenarios and numerous approaches have been adapted in the literature to modify spacetime metrics and eliminate CTCs. Some treatments suggest removing or modifying problematic regions of spacetime, such as cutting out those responsible for CTCs [20]. Also, Stephen Hawking proposed the ‘chronology protection conjecture’, hypothesizing that quantum effects prevent the formation of CTCs in physically realistic spacetimes [30]. Hawking famously said there seems to be a chronology protection agency that prevents the emergence of CTCs, thus making the universe safe for historians. Also, adjusting the boundary conditions at infinity is a way to avoid CTCs [29]. Nonetheless, such approaches to eliminate CTCs in the $4$-dimensional Taub-NUT metric come with trade-offs such as restricting spatial coordinates [20].
In this work, we investigate the Lorentzian Taub-NUT solution, retaining its CTCs and the associated causal structure. Taking advantage of the formal similarity of the Taub-NUT metric and the KK metric, we show that when the NUT charge is non-zero, time can be considered as an extra dimension that can be curled by reducing the $(1+3)$ dimensions to $3$-spatial dimensions. The outline of the work is as follows. In the next section, we briefly review the KK theory as the Lagrangian dimensional reduction procedure. In Sec. $3$, the Taub-NUT metric and the notion of the NUT charge are introduced. In Sec. $4$, we explain how the unification of gravitation and a gravitomagnetism, which originates from the NUT charge, takes place. Finally, we summarize the results obtained in the conclusions section. However, we include some derived relations in an appendix.
2 Brief Review of the Kaluza-Klein Theory
The KK theory approach to unification of electromagnetism and gravitation involves introducing a $5$-dimensional metric [^5] We represent $5$-dimensional quantities with a hat, $\hat{}$ , over them.
$(\hat{g}_{\rm AB})=\left(\begin{array}[]{cc}{g}_{\mu\nu}+\kappa^{2}A_{\mu}A_{\nu}\phi^{2}&\;\;\kappa A_{\mu}\phi^{2}\\ \kappa A_{\nu}\phi^{2}&\;\;\phi^{2}\end{array}\right),$ (1)
where $\phi$ is a scalar field, the uppercase Latin indices run from zero to four, and the lowercase Greek indices run from zero to three. Also, the metric tensor $g_{\mu\nu}$ is exactly the $4$-dimensional metric tensor in the Einstein theory of gravitation, the vector $A_{\mu}$ is the electromagnetic vector potential independent from the scalar field, $\kappa$ is a constant that can be expressed in terms of the $4$-dimensional Newtonian gravitational constant $G$ as
$\kappa\equiv 4\sqrt{\pi G}.$ (2)
Interestingly, a simple calculation shows that $\hat{g}=\phi^{2}\,g$, where $\hat{g}$ and $g$ are each the determinant of the corresponding metrics. The $4$-dimensional metric signature is taken to be (,+,+,+)(-,+,+,+), and we work in the natural units with $c=1=\hbar$, in which $G$ has the dimension of the square of length. Metric (1) can also be written as
$d{\hat{s}}^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}+\phi^{2}(dy+\kappa A_{\mu}dx^{\mu})^{2},$ (3)
where $(x^{\mu})=(x^{0},x^{1},x^{2},x^{3})$ are the usual four coordinates and $y=x^{4}$ is the fifth coordinate.
The $5$-dimensional analogue of the Einstein-Hilbert action in vacuum can be written as
$\hat{S}=\frac{1}{\hat{\kappa}^{2}}\int\hat{R}\sqrt{-\hat{g}}\,d^{4}x\,dy,$ (4)
where $\hat{R}$ is the $5$-dimensional Ricci scalar, and $\hat{\kappa}$ can be expressed in terms of a $5$-dimensional counterpart gravitational constant $\hat{G}$ with the dimension of length to the power of three in the natural units as $\hat{\kappa}\equiv 4\sqrt{\pi\hat{G}}$. By applying the cylinder condition, i.e. eliminating all derivatives with respect to the fifth coordinate, $\partial_{4}\,\hat{g}_{\rm AB}=0$, the $4$-dimensional effective action can be obtained through the dimensional reduction procedure as [13]
$S^{\rm(KK)}\equiv\hat{S}=\frac{1}{{\kappa^{2}}}\int d^{4}x\,\sqrt{-g}\,\phi\,\left(R-\frac{1}{4}\phi^{2}\kappa^{2}F_{\alpha\beta}F^{\alpha\beta}+2\,\frac{\partial^{\alpha}\phi\,\partial_{\alpha}\phi}{\phi^{2}}\right),$ (5)
where $R$ is the Ricci scalar, and $F_{\alpha\beta}\equiv\partial_{\alpha}A_{\beta}-\partial_{\beta}A_{\alpha}$ is the electromagnetic field tensor. Also, we have compactified the fifth dimension and have considered
$G=\frac{\hat{G}}{\oint dy}.$ (6)
Now, by varying the vacuum action (4) with respect to the $5$-dimensional metric $\hat{g}_{AB}$ and then performing the dimensional reduction procedure, one obtains the corresponding Einstein and Maxwell equations, and the modified Klein-Gordon equation [^6] Note that the metric signature used in this work is different from the one used in Ref. [13]. Additionally, see the appendix for some points on deriving field equations. [9, 13]
$G_{\mu\nu}\!\!\!$ $=$ $\!\!\!\frac{\kappa^{2}\phi^{2}}{2}T_{\mu\nu}^{\rm(EM)}+\frac{1}{\phi}\Big[\nabla_{\mu}(\partial_{\nu}\phi)-g_{\mu\nu}\Box\phi\Big],$
$\nabla^{\mu}\,F_{\mu\nu}\!\!\!$ $=$ $\!\!\!-3\,\frac{\partial^{\mu}\phi}{\phi}\,F_{\mu\nu}\qquad\textit{\rm and}\qquad\Box\phi=\frac{\kappa^{2}\phi^{3}}{4}\,F_{\mu\nu}F^{\mu\nu},$ (7)
where $G_{\mu\nu}=R_{\mu\nu}-Rg_{\mu\nu}/2$ is the $4$-dimensional Einstein tensor, $T_{\mu\nu}^{\rm(EM)}\equiv F_{\mu\alpha}F_{\nu}{}^{\alpha}-g_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}/4$ is the electromagnetic energy-momentum tensor, [^7] For a slowly varying scalar field, one has $\nabla_{\mu}\,F^{\mu\nu}=0$, which is the source-free Maxwell equations. and $\Box\equiv\nabla_{\mu}\nabla^{\mu}$.
When the scalar field $\phi$ is constant throughout spacetime, Eqs. (7) become
$G_{\mu\nu}=\frac{\kappa^{2}\phi^{2}}{2}T_{\mu\nu}^{\rm(EM)},\qquad\qquad\nabla^{\mu}\,F_{\mu\nu}=0\qquad\quad\textit{\rm and}\qquad\quad\Box\phi=0,$ (8)
where the first two equations are the corresponding Einstein and the homogeneous Maxwell equations, the third equation describes a massless Klein-Gordon scalar field known as the dilaton field. It is worthwhile to mention that the condition $\phi$ being a constant is valid only when $F_{\mu\nu}F^{\mu\nu}=0$ [36][38]. The main achievement of the KK theory is that starting with a source free action (4) leads to Eqs. (8) with a $4$-dimensional electromagnetic source matter originating purely from the geometry of $5$-dimensional empty spacetime.
3 The Taub-NUT Solution
The Taub-NUT solution is an important exact solution to the Einstein field equations. It represents a class of spacetimes that generalize the Schwarzschild solution by incorporating additional parameters associated with gravitational monopoles and dipole moments. The analogue of this solution in $5$-dimensional spacetime is called the Gross-Perry solution [39]. The Taub-NUT solution, which is a vacuum solution and exhibits unique topological and geometric properties, has been expressed in the spherical spatial coordinates as [20, 40]
$ds^{2}=-\Phi^{2}(r)\left(dt+2l\cos{\theta}d{\varphi}\right)^{2}+\frac{dr^{2}}{\Phi^{2}(r)}+(r^{2}+l^{2})\left(d\theta^{2}+\sin^{2}{\theta}d\varphi^{2}\right),$ (9)
where its metric function is obtained as
$\Phi(r)=\sqrt{\frac{r^{2}-2Mr-l^{2}}{r^{2}+l^{2}}},$ (10)
for when $r^{2}-2Mr-l^{2}>0$, which implies [^8] However, the range of $r$ in the spherical coordinates is $\geq 0. r<r$$r_{-}=M-\sqrt{M^{2}+l^{2}}$ and/or $r>r_{+}=M+\sqrt{M^{2}+l^{2}}$. Here, $M$ represents the geometric mass and $l$ the NUT charge, both of which have the dimension of length. The NUT charge $l$ distinguishes the Taub-NUT solution from the Schwarzschild metric, which corresponds to the case $l=0$. Such a NUT parameter introduces a non-trivial structure in spacetime, which is often interpreted as a gravitomagnetic monopole moment. The presence of the term $2l\cos{\theta}d{\varphi}$ in the metric indicates a coordinate singularity along the $z$-axis, known as the Misner string [29], which is a semi-infinite (quasi-regular) singularity along the half-axis $\theta=\pi$. Such a string is associated with non-trivial topology and is analogous to the Dirac string in the theory of magnetic monopoles [28]. Unlike the Schwarzschild solution, the Taub-NUT spacetime is not asymptotically flat in the usual sense and exhibits more complex asymptotic behavior due to the parameter $l$.
The main significance of the Taub-NUT solutions is the existence of the Misner string at $\theta=0$ or $\theta=\pi$, which makes the $1$-form $d\varphi$ ill-defined at either of these points. To get rid of the singularity at $\theta=0$, one can shift the time coordinate such that $t^{\prime}=t+2n\varphi$, and for the point $\theta=\pi$, one can take $t^{\prime\prime}=t-2n\varphi$, where $n$ is an integer. Since $\varphi$ is periodic by $2\pi$, one has to identify $t^{\prime}$ and $t^{\prime\prime}$ by modulo $8\pi n$, which boils down to imposing the periodicity to the time coordinate as $t\rightarrow t+8\pi n$. Doing this makes the $1$-form $d\varphi$ well-defined but leads to CTCs in the Lorentzian metric. Importantly, the elimination of the Misner string also leads to a quantization condition, namely the quantization of the NUT charge [41], which also acts as a topological charge by determining the topological twisting of spacetime [42].
4 Time as an Extra Dimension in Taub-NUT Metric
To address the CTC issue, in this section we try to establish a connection between the Taub-NUT solution and the KK framework. The main purpose is to start from the Taub-NUT metric in a $4$-dimensional spacetime coordinate, when $l\neq 0$, and eventually arrive at a $3$-dimensional spatial metric plus a compactified or curled time dimension. Actually, we assume that there is a dynamical mechanism that drags the size of time dimension down and keeps it in a stable way at an unobservably tiny scale. To begin, we rewrite the $4$-dimensional Taub-NUT metric (9) in the matrix form [^9] We represent $3$-dimensional quantities with a tilde, $\widetilde{}$ , over them.
$(g_{\mu\nu})=\left(\begin{array}[]{cc}-\Phi^{2}(r)&\;\;-2l\tilde{A}_{a}\Phi^{2}(r)\\ -2l\tilde{A}_{b}\Phi^{2}(r)&\;\;\tilde{g}_{ab}-4l^{2}\tilde{A}_{a}\tilde{A}_{b}\Phi^{2}(r)\end{array}\right),$ (11)
where the lowercase Latin indices run from one to three, the solution to the $3$-dimensional vector potential is obtained as $\tilde{\bf A}=(0,0,\cos{\theta})$, the metric signature is (,+,+,+)(-,+,+,+) provided that $\tilde{g}_{33}-4l^{2}\Phi^{2}(r)\cos^{2}{\theta}>0$, and $g=-\Phi^{2}(r)\,\tilde{g}$, where $\tilde{g}$ is the determinant of the $3$-dimensional metric. Moreover, the solution to the $3$-dimensional metric is
[…]
with the signature $(+,+,+)$ in the spherical coordinates.
Following the same steps as in Sec. (2), we proceed to write the $4$-dimensional spacetime action in vacuum as
${S}=\frac{1}{\kappa^{2}}\int{R}\sqrt{-g}\,d^{3}x\,dx^{0},$ (13)
where $x^{0}$ is the Killing vector of the Taub-NUT metric and can fulfil a compact direction, and $\kappa$ as defined in relation (2). Then, by applying the cylinder condition, i.e. eliminating all derivatives with respect to the $x^{0}$-coordinate, $\partial_{0}\,{g}_{\mu\nu}=0$, we can achieve the $3$-dimensional effective action from action (13) via the dimensional reduction procedure as [^10] The ansatz $\Psi^{2}\equiv-\Phi^{2}$ (and hence $\Psi=i\,\Phi$) simply makes metric (11) similar to metric (1) and in turn the same for the other relations.
$\tilde{S}\equiv S=\frac{1}{\tilde{\kappa}^{2}}\int d^{3}x\,\sqrt{\tilde{g}}\,\Phi(r)\,\left(\tilde{R}+\frac{1}{4}\Phi^{2}(r)\,4l^{2}\tilde{F}_{ab}\tilde{F}^{ab}-2\,\frac{\partial^{1}\Phi(r)\,\partial_{1}\Phi(r)}{\Phi^{2}(r)}\right),$ (14)
where $\tilde{R}$ is the $3$-dimensional Ricci scalar and $\tilde{F}_{ab}=\partial_{a}\tilde{A}_{b}-\partial_{b}\tilde{A}_{a}$ is an electromagnetic-like field tensor in three dimensions. Meanwhile, we have analogously introduced a constant $\tilde{\kappa}$ which can be expressed in terms of a $3$-dimensional counterpart gravitational constant $\tilde{G}$ with the dimension of length in the natural units as $\tilde{\kappa}\equiv 4\sqrt{\pi\tilde{G}}$. Actually, we have compactified the $x^{0}$-coordinate and have considered
$\tilde{G}\equiv\frac{{G}}{\oint dx^{0}},$ (15)
where $\oint dx^{0}$} is the result of integration along the $x^{0}$-direction. From the $3$-dimensional point of view, $\oint dx^{0}$} can be considered as the size of the CTC. The reduced action involves the gauge field $\tilde{A}_{a}$, which is a gravitomagnetic type gauge field associated with the NUT charge, and $\Phi(r)$, which plays the role of a scalar or a dilaton field within the context of the KK theory.
The $3$-dimensional equations of motion from the variation of the vacuum action (13) with respect to the $4$-dimensional metric $g_{\mu\nu}$ and then performing the dimensional reduction procedure, are [^11] Also, see Ref. [43].
$\tilde{G}_{ab}\!\!\!$ $=$ $\!\!\!-2l^{2}\Phi^{2}(r)\tilde{T}_{ab}^{(\rm GM)}+\frac{1}{\Phi(r)}\left[\tilde{\nabla}_{a}(\partial_{b}\Phi(r))-\tilde{g}_{ab}\tilde{\Box}\Phi(r)\right],$
$\tilde{\nabla}^{a}\,\tilde{F}_{ab}\!\!\!$ $=$ $\!\!\!-3\,\frac{\partial^{a}\Phi(r)}{\Phi(r)}\,\tilde{F}_{ab}\qquad{\rm and}\qquad\tilde{\Box}\Phi(r)=-l^{2}\Phi^{3}(r)\tilde{F}_{ab}\tilde{F}^{ab},$ (16)
where $\tilde{G}_{ab}$ is the $3$-dimensional Einstein tensor and $\tilde{T}_{ab}^{(\rm GM)}\equiv\tilde{F}_{ac}\tilde{F}_{b}{}^{c}-\tilde{g}_{ab}\tilde{F}_{cd}\tilde{F}^{cd}/4$ is a gravitomagnetic energy-momentum tensor, associated with the NUT charge, in three spatial dimensions whose trace is non-zero, in fact $\tilde{T}^{(\rm GM)}=\tilde{F}_{cd}\tilde{F}^{cd}/4$. Eqs. (16) represent the corresponding Einstein, gravitomagnetic and Klein-Gordon equations in three spatial dimensions, and in the appendix, we will mention some relations and calculations in this regard. Alternatively, utilizing some of the relations presented in the appendix, the third equation of Eqs. (16) gives the second-order differential equation
$\Phi^{2}\Phi^{\prime\prime}+\Phi(\Phi^{\prime})^{2}+\frac{2\,r\Phi^{2}\Phi^{\prime}}{r^{2}+l^{2}}+\frac{2l^{2}\Phi^{3}}{(r^{2}+l^{2})^{2}}=0,$ (17)
which is satisfied by the metric function (10) as its solution. Similarly, and also using $\tilde{\bf A}=(0,0,\cos{\theta})$, the $3$-dimensional metric (12) does satisfy the first equation in Eqs. (16).
Now, to see how the model deals with the CTC problem, we note that for $r_{-}<r<r_{+}$ including $r=l$, the scalar field $\Phi(r)$ (10) becomes imaginary and hence the signature of metric (9) changes to $(+,+,+,+)$. Such changes in signature indicate a shift from $(3+1)$ spacetime to a $4$-dimensional spatial geometry, or equivalently, a shift from the Lorentzian to the Euclidean geometry. Therefore, the CTC does not appear in the model in this range. As an extension of this work, the effects of such a non-constant scalar field can also be considered within the framework of the Brans-Dicke theories [44].
Our analysis indicates that when time is not compactified, the Taub-NUT metric unifies the gravitomagnetism and gravitation in $(1+3)$-dimensional spacetime. On the other hand, when the time coordinate is considered curled to a non-observable size, similar to the fifth dimension of the KK theory, gravitation and the gravitomagnetism, associated with the NUT charge, appear as two distinct fields. We emphasize that such a compactification does not work when only the ordinary mass is involved and the NUT charge is zero.
Furthermore, analogously comparing metric (11) with metric (1), action (14) with action (5), and Eqs. (16) with Eqs. (7), implies that the constant $\tilde{\kappa}$ is likely to be related to the NUT charge $l$. In this regard, considering their dimension (unit), we plausibly infer such a relation to be
$\tilde{\kappa}^{2}\longleftrightarrow 2l.$ (18)
Accordingly, utilizing the definition $\tilde{\kappa}$ in terms of $\tilde{G}$ into relation (18), while using relation (15), provides a relation between the $4$-dimensional gravitational constant and the NUT charge as
$G\longleftrightarrow\left(\frac{\oint dx^{0}}{8\pi}\right)l$ (19)
in the natural units. By analyzing the geometric properties, such a relation (19) between the NUT charge and the Newton gravitational constant may be of interest.
5 Conclusions
The KK theory unifies gravitation and electromagnetism by adding an extra dimension, while assuming that this extra dimension is curled to a very small scale, such as the Planck scale. After compactification in four dimensions, the KK metric leads to Einstein equations and Maxwell equations of electromagnetism. On the other hand, the Lorentzian Taub-NUT metric, as a peculiar solution to the Einstein field equations of GR in $(1+3)$ dimensions, introduces a parameter called the NUT charge that leads to a gravitomagnetic field, which originates from the gravitational magnetic monopole and does not arise in the weak-field slow-motion limit of GR. Such a monopole exhibits a feature called the Misner string, a singularity analogous to the Dirac string in electromagnetism. One way to avoid the Misner string is to make time periodic, which automatically leads to CTCs. However, the existence of CTCs challenges the causality principle and has led to the idea that Taub-NUT is more of a gravitational toy model than a realistic physical solution.
In this work, we investigate the Lorentzian Taub-NUT solution, retaining its CTCs and the associated causal structure. Then, taking advantage of the formal similarity of the Taub-NUT metric and the KK metric and when the NUT charge is non-zero (i.e., when it is not the Schwarzschild limit), we have considered time as an extra dimension in four dimensions. After such an extra dimension is curled, i.e. by reducing the $(1+3)$ dimensions to $(3)$ spatial dimensions, the Einstein vacuum equations in $(1+3)$ dimensions lead to the corresponding Einstein field equations in $(3)$ spatial dimensions plus the gravitomagnetic Maxwell-type equations related to the gravitational monopole. Such a gravitomagnetic field also acts as a source for the $3$-dimensional Einstein equations, similar to the KK reduction, in which the electromagnetic field acts as a source for $(1+3)$-dimensional Einstein equations.
In this process, we have attained gravitation and the gravitomagnetism, associated with the NUT charge, as two distinct fields. In fact, in this way, a new gravitomagnetic field has emerged from the temporal dimensional reduction. Moreover, such a compactification leads to a specific form of the dilaton field, which can have interesting effects in the framework of the Brans-Dicke theories. Remarkably, the CTC issue is resolved because for $r_{-}<r<r_{+}$ including $r=l$, the time direction becomes spacelike, changing the $(1+3)$ spacetime to a $4$-dimensional spatial geometry, and hence, the CTC does not appears in the model in this range.
We should point out that although the results have been derived from classical field equations in the Lorentzian signature, they may admit a Euclidean counterpart, subject to appropriate analytical continuation and boundary conditions, in which the NUT charge can be interpreted as a quantized topological charge, arising from the non-trivial fiber bundle structure and the regularity conditions in the Euclidean gravity [45].
Appendix
Action (5) as it stands, when $\phi$ is enclosed inside the parentheses, is actually a Barns-Dicke action. Hence, when varying it with respect to the $4$-dimensional metric $g_{\mu\nu}$, the scalar field $\phi$ and the vector potential $A_{\mu}$, the effect of its last (kinetic) term (i.e., $\partial^{\alpha}\phi\,\partial_{\alpha}\phi/\phi$) appears as the additional term $(g_{\alpha\beta}\partial^{\mu}\phi\,\partial_{\mu}\phi/2-\partial_{\alpha}\phi\,\partial_{\beta}\phi)/\phi^{2}$ in the first equation of Eqs. (7). Whereas, such a term does not exist there because action (5) is obtained by the dimensional reduction procedure of action (4). Indeed, for the last term inside the parentheses of action (5), one can use the relation $\partial^{\alpha}\phi\,\partial_{\alpha}\phi/\phi^{2}=\Box\phi/\phi-\nabla_{\mu}(\partial^{\mu}\phi/\phi)$. Since the second term in this relation is a complete derivative, when considered in action (4), it does not contribute in the corresponding field equations and can be ignored. Then, the remaining first term in this relation, when considered in action (5) and $\phi$ is placed inside the parentheses, is also a complete derivative and hence, does not contribute in the corresponding field equations either. Similar argument holds for action (14) and its corresponding field Eqs. (16).
Meanwhile, utilizing relations (10) and (12) and the definition of $\tilde{F}_{ab}$, and calculating the Ricci and Einstein tensors $\tilde{R}_{ab}$ and $\tilde{G}_{ab}$, and the Ricci scalar R, while using the Christoffel symbols provided in the following, it is straightforward to show that Eqs. (16) hold. For example, to prove the last equation of Eqs. (16), utilizing the standard relation
$\tilde{\Box}\Phi=\frac{1}{\sqrt{\mid\tilde{g}\mid}}\partial_{a}\left(\sqrt{\mid\tilde{g}\mid}\,\tilde{g}^{ab}\partial_{b}\Phi\right),$ (20)
its right hand-side, while using relations (10) and (12), gives
$\tilde{\Box}\Phi(r)=-\frac{2l^{2}\left(\frac{r^{2}-2Mr-l^{2}}{r^{2}+l^{2}}\right)^{\frac{3}{2}}}{(r^{2}+l^{2})^{2}}.$ (21)
On the other hand, using the definition of $\tilde{F}_{ab}$, its non-zero components are
$\tilde{F}_{23}=-\tilde{F}_{32}=-\sin{\theta},\qquad{\rm and\ in\ turn}\qquad\tilde{F}_{ab}\tilde{F}^{ab}=2\tilde{g}^{22}\tilde{g}^{33}\tilde{F}_{23}\tilde{F}_{23}=\frac{2}{(r^{2}+l^{2})^{2}}.$ (22)
Also, from relation (10), we obviously get
$\Phi^{3}(r)=\left(\frac{r^{2}-2Mr-l^{2}}{r^{2}+l^{2}}\right)^{\frac{3}{2}},$ (23)
hence via these two relations, the last equation of Eqs. (16) is verified.
Likewise, it is easy to prove the other two equations of Eqs. (16). However, in this regard, using the $3$-dimensional metric tensor (12) with function (10), the corresponding non-zero Christoffel symbols read
$\tilde{\Gamma}^{1}{}_{11}\!\!\!\!\!$ $=$ $\!\!\!\!\!\frac{Ml^{2}-2l^{2}r-Mr^{2}}{(r^{2}-2Mr-l^{2})(r^{2}+l^{2})},\qquad\qquad\tilde{\Gamma}^{1}{}_{22}=-\frac{r(r^{2}-2Mr-l^{2})}{r^{2}+l^{2}},$
$\tilde{\Gamma}^{1}{}_{33}\!\!\!\!\!$ $=$ $\!\!\!\!\!\tilde{\Gamma}^{1}{}_{22}\sin^{2}{\theta},\qquad\qquad\qquad\qquad\qquad\ \tilde{\Gamma}^{2}{}_{33}=-\cos{\theta}\sin{\theta},$
$\tilde{\Gamma}^{2}{}_{12}\!\!\!\!\!$ $=$ $\!\!\!\!\!\tilde{\Gamma}^{2}{}_{21}=\frac{r}{r^{2}+l^{2}}=\tilde{\Gamma}^{3}{}_{13}=\tilde{\Gamma}^{3}{}_{31},\qquad\tilde{\Gamma}^{3}{}_{23}=\tilde{\Gamma}^{3}{}_{32}=\cot{\theta}.$ (24)
Hence, the corresponding $3$-dimensional non-zero components of the Einstein tensor become
$\tilde{G}_{11}\!\!\!\!\!$ $=$ $\!\!\!\!\!-\frac{2Mr^{3}+3l^{2}r^{2}+l^{4}}{(r^{2}-2Mr-l^{2})(r^{2}+l^{2})^{2}},\qquad\tilde{G}_{22}=\frac{Mr^{3}+3l^{2}r^{2}-3Ml^{2}r-l^{4}}{(r^{2}+l^{2})^{2}},$ (25)
$\tilde{G}_{33}\!\!\!\!\!$ $=$ $\!\!\!\!\!\tilde{G}_{22}\sin^{2}{\theta}.$ (26)
Acknowledgements
The central concept and theoretical framework of this work was conceived and developed by M. Z.-Abyaneh, who also had the primary role in the implementation and authored the initial draft of the manuscript. M. Farhoudi had the primary role in editing and calculations, to which both authors contributed. Both authors engaged in constructive theoretical discussions throughout the project, and worked collaboratively to refine and finalize the manuscript. M. Z.-Abyaneh also would like to thank Prof. Ingemar Bengtsson for helpful discussions.